# PatentMind: A Multi-Aspect Reasoning Graph for Patent Similarity Evaluation

### v2

@@ 1 Introduction
Patent documents pose significant challenges for NLP-based similarity evaluation due to specialized domain knowledge, intricate legal language, and complex structural formats (Aslanyan and Wetherbee, 2022; Indukuri et al., 2007; Casola and Lavelli, 2021; Lupu and Hanbury, 2013). Accurate evaluation of patent similarity is crucial for identifying prior art, assessing infringement risks, and determining the novelty of innovations (Feng, 2020; Helmers et al., 2019; Arts et al., 2018; Wang et al., 2023). Given that inaccuracies can lead to overlooked prior art, invalid patent grants, or undetected infringements, robust and interpretable evaluation methods are essential. Moreover, the rapidly growing volume of patent applications across diverse technological fields further intensifies the need for reliable similarity evaluation techniques (Jiang and Goetz, 2024). Given the increasing reliance on patent analytics for R&D strategy, legal risk assessment, and competitive intelligence, developing accurate and interpretable similarity models has become critically important in both industry and academia (Hain et al., 2022; Yoo et al., 2023).
@@
In the early stages of patent similarity evaluation, keyword-based methods such as bag-of-words were widely used. However, these approaches struggled to capture the semantics and technical specificity of patent documents, resulting in ineffective representations (D’hondt et al., 2013; Ascione and Sterzi, 2024). To overcome these limitations, embedding-based methods such as static word embeddings and transformer-based contextual embeddings were subsequently introduced (Lee and Hsiang, 2020; Reimers and Gurevych, 2019). Nevertheless, these methods compress patent texts into dense vectors, which can obscure critical distinctions and reduce interpretability.
Recent advances in large language models (LLMs) have demonstrated considerable promise in reasoning-based tasks, offering a compelling alternative to traditional embedding-based approaches. Unlike fixed embedding methods, LLMs support dynamic reasoning and semantic comprehension, facilitating more sophisticated similarity assessments. However, directly applying general-purpose LLMs to patent similarity tasks remains challenging due to specialized terminology and complex features inherent in patent texts (Ikoma and Mitamura, 2025; Ascione and Sterzi, 2024). Moreover, prompt-based approaches to LLMs have partially addressed the limitations of embeddings, but still struggle due to the specialized terminology and structural complexity of patent documents.
To address these challenges, we propose PatentMind, a novel patent similarity evaluation framework structured as a Multi-Aspect Reasoning Graph (MARG). This framework emulates the analytical processes of patent experts through multi-step reasoning based on LLMs. PatentMind decomposes patent documents into three core dimensions: technical features, application domains, and claim scopes. Each dimension is independently evaluated for similarity, and the results are integrated through a context-aware reasoning process consisting of four stages: Domain Relationship Analysis, Information Distribution, Dimension Relevance, and Cross-Validation. This multi-dimensional approach is designed to closely reflect real-world patent examination practices. (USPTO, 2024).
@@ 2.1 Patent Similarity Evaluation
Early approaches to patent similarity evaluation, such as keyword matching and TF-IDF, treated patents as generic texts and often failed to capture their domain-specific structure and terminology (Tseng et al., 2007; Ascione and Sterzi, 2024). To address these limitations, vector-based models like Word2Vec and Doc2Vec were introduced to encode semantic relationships within patent corpora (Jeon et al., 2022). However, these models struggled with challenges such as polysemy and limited domain adaptation. To improve semantic representation, contextualized language models like SciBERT (Beltagy et al., 2019) and PatentBERT (Lee and Hsiang, 2020) were developed, offering embeddings tailored for scientific and patent texts. While these models produce more accurate representations, their black-box nature and lack of interpretability hinder their practical application in legal and technical decision-making contexts (Miric et al., 2023; Ascione and Sterzi, 2024).
@@ 2.2 Prompt Engineering
Prompt engineering has emerged as a key strategy for enabling complex reasoning in LLMs. Chain-of-Thought prompting (Wei et al., 2022) improves sequential reasoning by guiding models through intermediate steps. However, its linear structure limits effectiveness in tasks requiring divergent reasoning. To overcome this, Yao et al. (Yao et al., 2023) proposed the Tree-of-Thoughts framework, which enables exploration of multiple reasoning paths in parallel, enhancing performance on tasks requiring structured deliberation. Additional work demonstrates that carefully designed prompts enable strong few-shot (Brown et al., 2020) and zero-shot (Kojima et al., 2022) reasoning, using minimal task examples or simple cues. These prompting strategies significantly improve LLM performance on complex tasks, laying the foundation for PatentMind’s structured reasoning approach, where multi-dimensional analysis is essential for capturing both technical and legal dimensions of patents.
@@ 3.1 Multi-Dimensional Feature Extraction
PatentMind captures the complex structure and domain-specific nuances of patent documents through LLM-based feature extraction. We decompose each patent into three dimensions: technical feature ($T$), application domain ($D$), and claim scope ($C$). These dimensions reflect how patent specialists evaluate patent similarity, focusing on what the invention is, where it applies, and how broadly it is claimed (USPTO, 2024).
To obtain these dimensions, we apply an LLM-based feature extraction function, $T$, $D$, or $C$, on a patent document $P$ to obtain corresponding content. The function generates a structured representation composed of technical features $T(P)$, application domains $D(P)$, and claim scope $C(P)$. The title, abstract, and claims of the patent are used as input for the extractor. This structured prompting approach guides the LLM using aspect-specific instructions. Each prompt focuses on a distinct semantic aspect: methodologies for technical features, application contexts for domains, and legal boundaries for claim scope. These prompts are grounded in structured inputs from the title, abstract, and claims. Prompt designs for all dimensions are provided in Appendix B.
@@ 3.3.2 Multi-perspective Analysis
The LLM performs multi-dimensional reasoning to guide the dynamic weighting process. Four reasoning functions are applied, each using specific inputs to generate signals for assigning weights to the similarity scores $S_{T}$, $S_{D}$, $S_{C}$.
  • Domain Relationship Analysis: In this step, the function $f_{rel}$ leverages the IPC codes of the two patents to assess their technical domain relationship using an LLM-based reasoning process. The output is a categorical label, $R_{domain}$, which determines the extent to which the domain similarity score ($S_{D}$) should influence the final similarity calculation. Detailed prompting instructions for this module are provided in Appendix D.
  • Information Distribution Analysis: This step uses the function $f_{dist}$ to examine the distribution pattern of the aspect-wise similarity scores $S_{T}$, $S_{D}$, and $S_{C}$. The output, $D_{info}$, indicates which dimensions contribute most significantly to the perceived similarity between the two patents. The corresponding prompts used in this reasoning process are provided in Appendix D.
  • Dimension Relevance Assessment: In this step, the function $f_{assess}$ evaluates the relative importance of each similarity dimension based on the content of the patent pair ($C_{pair}$) and their domain relationship label ($R_{domain}$). The output, $A_{rel}$, consists of dimension-wise relevance scores that guide the prioritization of weights in the final similarity computation. Detailed prompting instructions for this assessment module are provided in Appendix D.
  • Cross-validation Reasoning: This step uses the function $f_{valid}$ to assess the consistency between the predicted importance of each similarity dimension ($A_{rel}$) and the actual distribution of similarity scores ($D_{info}$), taking into account the full context of the patent pair ($C_{pair}$). If the most relevant dimension identified in $A_{rel}$ aligns with the dominant signal in $D_{info}$, the configuration is considered robust. The output, $V_{cross}$, is a continuous robustness score ranging from [0,1], serving as a reliability indicator for the current weighting scheme. The prompting details for this reasoning step are provided in Appendix D.
These reasoning steps provide a logical foundation for computing dynamic weights, ensuring that the final similarity score reflects both the intrinsic properties of the patent pair and their contextual characteristics. To formalize this process, we represent the multi-stage reasoning pipeline as a directed acyclic graph $G=(V,E)$. Each node $v_{i}\in V$ corresponds to one of four reasoning steps–Domain Relationship Analysis, Information Distribution Analysis, Dimension Relevance Assessment, and Cross-validation Reasoning. The edges $e_{ij}\in E$ represent directed dependencies between these steps, facilitating systematic transmission of contextual signals such as IPC similarity, score patterns, and alignment strength. This modular graph structure supports a message-passing mechanism that integrates interdependent reasoning outcomes into the final weight derivation.
@@ 4 Dataset Construction
To facilitate rigorous evaluation of patent similarity methods, we introduce PatentSimBench, the first expert-annotated benchmark dataset tailored for patent similarity tasks. PatentSimBench provides rationale-supported similarity scores aligned with real-world legal and technical judgments, making it particularly suitable for evaluating multi-step reasoning models. The dataset consists of 500 patent pairs sampled from the USPTO database (USPTO, 2025), with a balanced representation across all IPC sections. The International Patent Classification (IPC) categorizes patents into eight top-level domains, such as human necessities, chemistry, physics, and digital technologies (WIPO, 1975), ensuring comprehensive technical coverage across diverse fields. Similarity annotations are based on a 5-point Likert scale, with scores ranging from 1 to 5. Our dataset is publicly available on GitHub(It will be available upon accept).
@@ 4.1 Dataset Composition
PatentSimBench includes patent pairs evenly distributed across major technological domains, covering all IPC sections. Table 1 summarizes key dataset statistics. The high inter-annotator agreement scores, including Fleiss’ Kappa (0.588) and Cronbach’s Alpha (0.967), demonstrate robust consensus among expert annotators despite the inherent subjectivity in assessing patent similarity.
@@ 4.2 Annotation Process
Four domain experts with both legal and technical backgrounds annotated PatentSimBench using a 5-point Likert scale. The annotation guidelines were aligned with procedures commonly followed by patent examiners for similarity and prior art evaluation. The annotation process followed a two-phase protocol: (1) independent assessments by each expert, and (2) consensus refinement through structured discussions of divergent cases, during which annotators reviewed and adjusted their ratings based on shared rationales. For quality control, we excluded patent pairs with a post-discussion standard deviation greater than 2 ($\sigma_{\text{ratings}}>2$), ensuring high reliability in the final dataset. This process was guided by the predefined annotation guidelines described in Appendix F.
@@ 5 Experimental Results
We conducted comprehensive experiments on the PatentSimBench dataset to evaluate PatentMind against embedding-based models, prompt engineering strategies, and regression baselines. To assess the robustness and contribution of individual components, we performed ablation studies and cross-model evaluations using various LLMs. All experiments were carried out with standard evaluation metrics, using GPT-4o-mini as the backbone model with a fixed temperature of 0.2.
@@ 7.3 High-Error Case Analysis
We further analyze the 50 patent pairs with the highest prediction errors to identify potential systematic limitations. We observe four recurring patterns: (1) underestimation of highly similar patents (e.g., predicted 0.53 vs. actual 0.833 in metallurgy); (2) overestimation of dissimilar patents (e.g., predicted 0.50 vs. actual 0.250 in marking tech); (3) difficulty distinguishing functional differences despite textual overlap (e.g., 0.48 vs. 0.750 in mechanical parts); (4) challenges in interpreting domain terms, particularly in biotech (e.g., 0.24 vs. 0.500). These findings suggest that, while PatentMind effectively captures general semantic similarity, performance may be further improved by incorporating more explicit representations of functional intent and domain-specific vocabulary. Additional examples and discussion are provided in Appendix G.
@@ 8 Conclusion
In this paper, we proposed PatentMind, a novel framework for patent similarity evaluation that integrates linguistic, legal, and technical reasoning.
First, we introduced a multi-dimensional analysis that, for the first time, jointly considers the three core aspects of patents: technical features, application domains, and claim scope. This approach effectively captures the structural complexity of patent documents.
Second, we proposed context-aware weighting, where LLM-based reasoning dynamically adjusts the relative importance of each dimension depending on domain- and case-specific contexts. This is the first method to discriminatively capture key factors under varying patent contexts.
Third, we achieved expert-level agreement and demonstrated clear superiority over embedding-based and prompting-based baselines, showing that our approach goes beyond simple LLM calls or prompting strategies.
Fourth, we showed that PatentMind outperforms specialized patent-tuned models, demonstrating higher accuracy while being more efficient than fine-tuning approaches that require substantial time and computational resources.
Fifth, we established the model-agnostic robustness of PatentMind. It consistently delivers strong performance across diverse LLMs, ensuring generalizability and reproducibility without dependence on any single vendor or architecture.
Finally, we introduced PatentSimBench, the first expert-annotated benchmark for patent similarity, which provides a valuable resource for both academia and industry.
Taken together, these contributions extend beyond computational linguistics, offering a structured and semantically grounded foundation for addressing real-world decision-making tasks such as prior art search, infringement risk assessment, and novelty evaluation.
@@ Appendix C Appendix C: Prompts for Similarity Computation
Prompt : In this appendix, we provide the prompts used to guide the Large Language Model (LLM) in computing the similarity scores for each of the three dimensions: Technical Features ($S_{T}$), Application Domains ($S_{D}$), and Claim Scope ($S_{C}$). These prompts enable the LLM to assess the degree of overlap between two patent documents based on their extracted features.
@@ F.7 Quality Control Process
Appendix G Appendix G : Detailed High-Error Case Analysis

### v3

@@ 1 Introduction
Patent documents pose significant challenges for NLP-based similarity evaluation due to specialized domain knowledge, intricate legal language, and complex structural formats (aslanyan2022patents; indukuri2007similarity; casola2021summarization; lupu2013patent). Accurate evaluation of patent similarity is crucial for identifying prior art, assessing infringement risks, and determining the novelty of innovations (Feng2020; Helmers2019; arts2018text; Wang2023). Given that inaccuracies can lead to overlooked prior art, invalid patent grants, or undetected infringements, robust and interpretable evaluation methods are essential. Moreover, the rapidly growing volume of patent applications across diverse technological fields further intensifies the need for reliable similarity evaluation techniques (jiang2024nlp). Given the increasing reliance on patent analytics for R&D strategy, legal risk assessment, and competitive intelligence, developing accurate and interpretable similarity models has become critically important in both industry and academia (hain2022similarity; yoo2023novel).
@@
In the early stages of patent similarity evaluation, keyword-based methods such as bag-of-words were widely used. However, these approaches struggled to capture the semantics and technical specificity of patent documents, resulting in ineffective representations (D'hondt2013; ascione2024comparative). To overcome these limitations, embedding-based methods such as static word embeddings and transformer-based contextual embeddings were subsequently introduced (lee2020patentbert; reimers2019sentence). Nevertheless, these methods compress patent texts into dense vectors, which can obscure critical distinctions and reduce interpretability.
Recent advances in large language models (LLMs) have demonstrated considerable promise in reasoning-based tasks, offering a compelling alternative to traditional embedding-based approaches. Unlike fixed embedding methods, LLMs support dynamic reasoning and semantic comprehension, facilitating more sophisticated similarity assessments. However, directly applying general-purpose LLMs to patent similarity tasks remains challenging due to specialized terminology and complex features inherent in patent texts (ikoma2025can; ascione2024comparative). Moreover, prompt-based approaches to LLMs have partially addressed the limitations of embeddings, but still struggle due to the specialized terminology and structural complexity of patent documents.
To address these challenges, we propose PatentMind, a novel patent similarity evaluation framework structured as a Multi-Aspect Reasoning Graph (MARG). This framework emulates the analytical processes of patent experts through multi-step reasoning based on LLMs. PatentMind decomposes patent documents into three core dimensions: technical features, application domains, and claim scopes. Each dimension is independently evaluated for similarity, and the results are integrated through a context-aware reasoning process consisting of four stages: Domain Relationship Analysis, Information Distribution, Dimension Relevance, and Cross-Validation. This multi-dimensional approach is designed to closely reflect real-world patent examination practices. (uspto_mpep_2141).
@@ 2.1 Patent Similarity Evaluation
Early approaches to patent similarity evaluation, such as keyword matching and TF-IDF, treated patents as generic texts and often failed to capture their domain-specific structure and terminology (tseng2007text; ascione2024comparative). To address these limitations, vector-based models like Word2Vec and Doc2Vec were introduced to encode semantic relationships within patent corpora (jeon2022doc2vec). However, these models struggled with challenges such as polysemy and limited domain adaptation. To improve semantic representation, contextualized language models like SciBERT (beltagy2019scibert) and PatentBERT (lee2020patentbert) were developed, offering embeddings tailored for scientific and patent texts. While these models produce more accurate representations, their black-box nature and lack of interpretability hinder their practical application in legal and technical decision-making contexts (miric2023using; ascione2024comparative).
@@ 2.2 Prompt Engineering
Prompt engineering has emerged as a key strategy for enabling complex reasoning in LLMs. Chain-of-Thought prompting (wei2022chain) improves sequential reasoning by guiding models through intermediate steps. However, its linear structure limits effectiveness in tasks requiring divergent reasoning. To overcome this, Yao et al. (yao2023tree) proposed the Tree-of-Thoughts framework, which enables exploration of multiple reasoning paths in parallel, enhancing performance on tasks requiring structured deliberation. Additional work demonstrates that carefully designed prompts enable strong few-shot (brown2020language) and zero-shot (kojima2022large) reasoning, using minimal task examples or simple cues. These prompting strategies significantly improve LLM performance on complex tasks, laying the foundation for PatentMind’s structured reasoning approach, where multi-dimensional analysis is essential for capturing both technical and legal dimensions of patents.
@@ 3.1 Multi-Dimensional Feature Extraction
PatentMind captures the complex structure and domain-specific nuances of patent documents through LLM-based feature extraction. We decompose each patent into three dimensions: technical feature ($T$), application domain ($D$), and claim scope ($C$). These dimensions reflect how patent specialists evaluate patent similarity, focusing on what the invention is, where it applies, and how broadly it is claimed (uspto_mpep_2141).
To obtain these dimensions, we apply an LLM-based feature extraction function, $T$, $D$, or $C$, on a patent document $P$ to obtain corresponding content. The function generates a structured representation composed of technical features $T(P)$, application domains $D(P)$, and claim scope $C(P)$. The title, abstract, and claims of the patent are used as input for the extractor. Note that these sections are legally mandated components for any valid patent application (e.g., 35 U.S.C. § 112), ensuring the availability of structured inputs. This structured prompting approach guides the LLM using aspect-specific instructions. Each prompt focuses on a distinct semantic aspect: methodologies for technical features, application contexts for domains, and legal boundaries for claim scope. These prompts are grounded in structured inputs from the title, abstract, and claims. Prompt designs for all dimensions are provided in Appendix B.
@@ 3.3.2 Multi-perspective Analysis
The LLM performs multi-dimensional reasoning to guide the dynamic weighting process through four interdependent functions (detailed prompts in Appendix D). First, Domain Relationship Analysis ($f_{rel}$) leverages IPC codes to output a categorical label $R_{domain}$, determining the influence of domain similarity. Second, Information Distribution Analysis ($f_{dist}$) examines the patterns of aspect-wise scores ($S_{T},S_{D},S_{C}$) to generate $D_{info}$. Third, Dimension Relevance Assessment ($f_{assess}$) evaluates dimension importance ($A_{rel}$) based on the full context $C_{pair}$ and $R_{domain}$. Finally, Cross-validation Reasoning ($f_{valid}$) checks the consistency between predicted relevance ($A_{rel}$) and actual score distribution ($D_{info}$), yielding a robustness score $V_{cross}\in[0,1]$. We formalize this pipeline as a directed acyclic graph $G=(V,E)$, where nodes correspond to reasoning steps and edges represent dependencies facilitating the message-passing of contextual signals for the final weight derivation.
@@ 4 Dataset Construction
To facilitate rigorous evaluation of patent similarity methods, we introduce PatentSimBench, the first expert-annotated benchmark dataset tailored for patent similarity tasks. PatentSimBench provides rationale-supported similarity scores aligned with real-world legal and technical judgments, making it particularly suitable for evaluating multi-step reasoning models. The dataset consists of 500 patent pairs sampled from the USPTO database (uspto_bulk_data), with a balanced representation across all IPC sections. The International Patent Classification (IPC) categorizes patents into eight top-level domains, such as human necessities, chemistry, physics, and digital technologies (wipo_ipc), ensuring comprehensive technical coverage across diverse fields. Similarity annotations are based on a 5-point Likert scale, with scores ranging from 1 to 5. Our dataset is publicly available on GitHub(It will be available upon accept).
@@ 4.1 Dataset Composition
PatentSimBench includes patent pairs evenly distributed across major technological domains, covering all IPC sections. Table 1 summarizes key dataset statistics. The moderate inter-annotator agreement scores, including Fleiss’ Kappa (0.588) and Cronbach’s Alpha (0.967), demonstrate robust consensus among expert annotators despite the inherent subjectivity in assessing patent similarity.
@@ 4.2 Annotation Process
Four domain experts with both legal and technical backgrounds annotated PatentSimBench using a 5-point Likert scale. The annotation guidelines were rigorously designed based on the USPTO Manual of Patent Examining Procedure (MPEP) §2141, which mandates a structured analysis of claim scope, technical content, and field of search. This ensures that our similarity criteria reflect actual legal standards rather than subjective judgments. The annotation process followed a two-phase protocol: (1) independent assessments by each expert, and (2) consensus refinement through structured discussions of divergent cases, during which annotators reviewed and adjusted their ratings based on shared rationales. For quality control, we excluded patent pairs with a post-discussion standard deviation greater than 2 ($\sigma_{\text{ratings}}>2$), ensuring high reliability in the final dataset. This process was guided by the predefined annotation guidelines described in Appendix F.
@@ 5 Experimental Results
We conducted comprehensive experiments on the PatentSimBench dataset to evaluate PatentMind against embedding-based models, prompt engineering strategies, and regression baselines. To assess the robustness and contribution of individual components, we performed ablation studies and cross-model evaluations using various LLMs. All experiments were carried out with standard evaluation metrics, using GPT-4o-mini as the backbone model. To ensure determinism and reproducibility, we used a fixed temperature of 0.2 and default sampling parameters (top_p=1.0). Furthermore, to verify result stability, we conducted 5 independent runs with different random seeds on a representative subset of the data. The observed standard deviation in Pearson correlation was negligible ($\sigma<0.001$), confirming the robustness of our results without the need for reporting averaged metrics over multiple trials.
@@ 5.4 Comparison of Regression Weighting
5.5 Efficiency and Cost-Benefit Analysis
To assess the practical viability of PatentMind, we evaluated its computational efficiency using GPT-4o-mini. On average, processing a single patent pair requires approximately 13,225 input tokens and 832 output tokens across 8 API calls. Through parallelized execution of independent extraction and scoring steps, the end-to-end latency is reduced to 20–30 seconds per pair. Financially, the operational cost is approximately \$0.003 USD per pair based on current API pricing. While computationally more expensive than simple embedding-based retrieval, this cost is negligible compared to human patent experts. As shown in Table 6, PatentMind offers expert-level reasoning at a fraction of the cost (<<\$0.01), making it a cost-effective solution for high-stakes decision-making tasks such as infringement risk assessment.
Table 6: Efficiency and cost comparison. Human cost is estimated based on average patent attorney hourly rates assuming 60 minutes per case.
Method Latency Cost / Pair
Embedding (SciBERT) << 0.1s $\approx$ \$0.000
PatentMind (Ours) 20 30s $\approx$ \$0.003
Human Expert >> 60 min >> \$100.0
We emphasize that this computational overhead is an intentional design choice. PatentMind is explicitly designed to function as a re-ranking and precision analysis module, typically applied after a broad initial retrieval phase. In high-stakes domains like patent infringement analysis or invalidity searches, where missing a critical prior art can result in significant legal consequences, accuracy and interpretability take precedence over millisecond latency. Therefore, the trade-off of higher computational cost for expert-level reasoning depth is justified for these targeted downstream applications.
@@ 7.3 High-Error Case Analysis
We further analyze the 50 patent pairs with the highest prediction errors to identify potential systematic limitations. We observe four recurring patterns: (1) underestimation of highly similar patents (e.g., predicted 0.53 vs. actual 0.833 in metallurgy); (2) overestimation of dissimilar patents (e.g., predicted 0.50 vs. actual 0.250 in marking tech); (3) difficulty distinguishing functional differences despite textual overlap (e.g., 0.48 vs. 0.750 in mechanical parts); (4) challenges in interpreting domain terms, particularly in biotech (e.g., 0.24 vs. 0.500). These findings suggest that, while PatentMind effectively captures general semantic similarity, performance may be further improved by incorporating more explicit representations of functional intent and domain-specific vocabulary. Additional examples and discussion are provided in Appendix H.
@@ 8 Conclusion
We proposed PatentMind, a novel framework integrating linguistic, legal, and technical reasoning for patent similarity. By jointly modeling technical features, application domains, and claim scope, it effectively captures structural complexity. A key innovation is our context-aware weighting mechanism, which uses LLM reasoning to dynamically adjust dimensional importance.
PatentMind achieves expert-level agreement ($r=0.938$) and model-agnostic robustness, significantly outperforming embedding, prompting, and patent-tuned baselines. Notably, it delivers higher accuracy with greater efficiency than resource-intensive fine-tuning approaches.
Finally, we release PatentSimBench, the first expert-annotated similarity benchmark. These contributions extend beyond computational linguistics, offering a semantically grounded foundation for critical real-world tasks such as prior art search, infringement assessment, and novelty evaluation.
@@ Appendix C Appendix C: Prompts for Similarity Computation
In this appendix, we provide the prompts used to guide the Large Language Model (LLM) in computing the similarity scores for each of the three dimensions: Technical Features ($S_{T}$), Application Domains ($S_{D}$), and Claim Scope ($S_{C}$). These prompts enable the LLM to assess the degree of overlap between two patent documents based on their extracted features.
@@ F.7 Quality Control Process
Appendix G Appendix G: Prompts for Baseline Methods
To ensure fair comparison and reproducibility, we provide the exact prompts used for the baseline methods, including Chain-of-Thought (CoT) and Few-shot prompting.
G.1 Prompt for Chain-of-Thought (CoT)
The CoT prompt encourages the model to generate intermediate reasoning steps before predicting the final score.
Prompt: You are an expert patent analyst. Your task is to evaluate the similarity between two patent documents, Patent A and Patent B.
Input:
- Patent A: [Title, Abstract, Claims]
- Patent B: [Title, Abstract, Claims]
Instructions:
1. Analyze the technical field, core innovation, and claim scope of both patents. 2. Compare the similarities and differences step-by-step. 3. Based on the reasoning, determine a similarity score between 0.0 (completely different) and 1.0 (identical).
Output the result in the following format:
Reasoning: [Your step-by-step analysis]
Score: [numerical score]
G.2 Prompt for Few-shot Prompting
For Few-shot prompting, we utilized a 3-shot setting. We selected three representative patent pairs with expert-annotated scores (Low, Medium, High) to guide the model.
Prompt: Evaluate the similarity between two patents on a scale of 0.0 to 1.0. Here are three examples:
Example 1:
Patent A: [Content of Patent A1]
Patent B: [Content of Patent B1]
Score: 0.2 (Low Similarity)
Example 2:
Patent A: [Content of Patent A2]
Patent B: [Content of Patent B2]
Score: 0.5 (Medium Similarity)
Example 3:
Patent A: [Content of Patent A3]
Patent B: [Content of Patent B3]
Score: 0.9 (High Similarity)
Task:
Patent A: [Target Patent A Content]
Patent B: [Target Patent B Content]
Score:
Appendix H Appendix H : Detailed High-Error Case Analysis